Kandi has $2.65 in nickels and dimes. If she has 40 coins in all, how many nickels does she have?
Answer by checkley71(8403)

265=10d+5n & d+n=40 or d=40-n then 265=10(40-n)+5n or 265=400-10n+5n or
264-400=-5n or 135=5n or n=27 number of nickels the 40-27=13 dimes
proof 10*13+5*27=265 or 130+135=265 or 265=265